### v3

# Interplay of orbital-selective Mott criticality and flat-band physics in La3Ni2O6

Abstract
Superconductivity in nickelates apparently takes place in two different Ni oxidation regimes, namely either for infinite-layer-type compounds close to Ni+, or for Ruddlesden-Popper materials close to Ni2+. The reduced La3Ni2O6 bilayer with a nominal Ni1.5+ oxidation state may therefore serve as a normal-state mediator between the two known families of $3d^{8}$-like and $3d^{9}$-like superconducting nickelates. Using first-principles many-body theory, we explain its experimental 55 meV charge gap as originating from a new type of correlated (quasi-)insulator. Flat-band electrons of Ni-$d_{z^{2}}$ character become localized from scattering with orbital-selective Mott-localized Ni-$d_{x^{2}-y^{2}}$ electrons, by trading in residual hopping energy for a gain in local exchange energy in a ferromagnetic Kondo-lattice scenario. Most importantly, the flat-band electrons offer another route to unconventional superconductivity in nickelates at ambient pressure.
[…]
In this context, the La3Ni2O6 compound (see Fig. 1a) stands out and may attain a unique role in connecting both superconducting nickelate families. As the fully reduced form of La3Ni2O7, it shares the bilayer structure but also displays the missing apical oxygens of the original reduced systems. Furthermore, its nominal Ni1.5+ oxidation state with $3d^{8.5}$ electron count just mediates between the latter $3d^{9}$-like systems and the ligand-hole affected $3d^{8}$-like RP compounds. Last but not least, it crystallizes in tetragonal symmetry, which is the common symmetry for superconducting nickelates.
Experimentally, the material has first been reported in 2006 in powder form [37], and soon afterwards characterized as semiconducting with no magnetic order down to 4 K [38]. Absence of both, static order and metallicity, has also been concluded from nuclear magnetic resonance experiments [39]. One may attribute semiconductivity to the polycrystalline samples, yet recent single-crystal studies confirmed non-metallicity, revealing a small activation gap of 55 meV at low $T$ [40]. Magnetization measurements show a weak kink at $T~\sim 176$ K, but otherwise again no obvious signature of a magnetic transition.
Figure 1: Reduced bilayer compound La3Ni2O6. (a) Tetragonal $I4/mmm$ crystal structure with La(green), Ni(blue) and O(red) atoms. (b) DFT dispersions in fatspec representation highlighting the Ni(3d)(3d) orbital characters.
[…]
Figure 2: Spectral properties of La3Ni2O6 from DFT+sicDMFT at $T=50$ K. (a) Upper panel: total and projected k-integrated spectral function (inset: low-energy blow up). Lower panel: Ni(3d)(3d) local spectral function. Right inset: low-energy blow up. (b) k-resolved spectral function $A({\bf k},\omega)$ along high-symmetry lines. (c) Fatspec representation of $A({\bf k},\omega)$ for the different Ni(3d)(3d) contributions (DFT bands: grey lines).
In this work, we report a theoretical description of the correlated electronic structure of paramagnetic La3Ni2O6 based on first-principles many-body theory. The experimental 55 meV gap is well reproduced, being the result of a novel correlated insulating scenario. It builds up on orbital-selective Mott physics [43, 44] of localized Ni-$d_{x^{2}-y^{2}}$ electrons acting on a Ni-$d_{z^{2}}$ flat-band feature. Interestingly, the flat-band low-energy dispersion also becomes gapped and the emerging semiconducting state is rather robust against hole doping $x$. It is only metallized for $x>0.15$ when the flat-band electrons are released from (quasi-)localization. The strong flat-band response at low-energy with threshold hole doping may signal further electronic instabilities including unconventional superconductivity with sign-changing $s_{\pm}$-wave symmetry.
Correlated electrons.— Previously, a realistic dynamical mean-field theory (DMFT) study of La3Ni2O7 [20] revealed half-filled Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals, respectively. Low-energy electrons in both orbitals are subject to strong correlations, yet the Ni-$d_{z^{2}}$ dispersion has substantial flat-band character at the Fermi level. As a result, the highly correlated metallic state is prone to instabilities, including superconductivity at high pressure. For electrons in the reduced bilayer La3Ni2O6 we also expect strong correlations, however, with much larger orbital differentiation due to the missing apical oxygens. As shown for infinite-layer NdNiO2 using the same theoretical framework [6], Ni-$d_{x^{2}-y^{2}}$ is even close to Mott localization, whereas “free-standing” Ni-$d_{z^{2}}$ turns out weakly correlated. Yet from formal $d^{9}$ in NdNiO2 the Ni-$d_{z^{2}}$ electron count is near complete filling, whereas from formal $d^{8.5}$ in La3Ni2O6 a much lower filling is expected.
The La3Ni2O6 inspection on the DFT level (see Fig. 1b) exhibits flat Ni-$d_{z^{2}}$ dispersions just below the Fermi level. But there is additional prominent low-energy flat-band character close to the X point from dominant Ni-$d_{xz,yz}$ character [45], as also observed for Nd3Ni2O6 [46]. To assess the role of correlations in the reduced bilayer, we discuss its electronic structure as described by the combination of DFT, self-interaction correction (SIC) and DMFT, i.e. within the DFT+sicDMFT approach [47] (see Supplemental Material with Refs. 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60 for further details). In the upper panel of Fig. 2a the total k integrated spectral function and the projected Ni(3d)(3d) and O(2p)(2p) content is displayed. A sizable charge-transfer energy $\Delta_{\rm ct}=\varepsilon_{d}-\varepsilon_{p}$, for respective band centers $\varepsilon_{d,p}$ of Ni(3d)(3d) and O(2p)(2p), leads to a spectral separation of both dominant chemical contributions. A broad O(2p)(2p) weight is centered around 4-4 eV, and the Ni(3d)(3d) weight dominantly peaks at $\sim-0.8$ eV. The latter peak is primarily formed by Ni-$t_{2g}$ character, as revealed by the local spectral function in the lower panel of Fig. 2a. Spectral integration leads to occupations $n_{p}=5.6$ and $n_{d}=8.3$, i.e. both lower than expected from nominal O2- and Ni1.5+. Hence there is a sizable ligand-hole character on oxygen, and a Ni(3d)(3d) count closer to the favorable $d^{8}$ occupation. This means that either the Ni(4s)(4s) or the La(5d,6s)(5d,6s) sub-shells are not fully oxidized. Note that while the total filling is physically well-defined, orbital fillings always depend on the choice of local functions. The substantial ligand-hole content for a nickelate with nominal $n_{d}>8$ is unexpected, though note that La3Ni2O6 is formally also located in an unusual $\Delta_{\rm ct}$ regime [61]. The obtained fillings still do not come as a total surprise, as an electride-like setting has e.g been put forward for infinite-layer nickelates [18].
Figure 3: Temperature-dependent Ni-$e_{g}$ data of La3Ni2O6 from DFT+sicDMFT. Left: Low-energy spectral function, and right: respective imaginary part of the Matsubara self-energy $\Sigma(i\omega_{n})$.
Figure 4: Spectral properties of hypothetical (La1-xSrx)3Ni2O6 at $T=50$ K based on DFT+sicDMFT using VCA for the doped La sites. (a) Total spectral function (inset: low-energy blow up). (b) Bloch-resolved spectral function $A_{\nu}(\omega)$ for $\nu=\nu_{\rm f}$ (variants of red) and for $\nu=\nu_{\rm f}\pm 1$ (grey), with $\nu_{\rm f}$ denoting the Bloch state, largest susceptible to hole doping. (c) Ni-$e_{g}$-resolved local spectral function (inset: low-energy blow up for Ni-$d_{z^{2}}$). (d,e) Fatspec representation of the k-resolved spectral function $A({\bf k},\omega)$, and the Fermi surface in the $k_{z}=0$ plane for $x=0.225$. Color coding of (d,e) as in Fig. 2c.
At first glance, the low-energy spectral weight at $T=50$ K seemingly shows a sharp quasiparticle(QP)-like peak, but the detailed resolution (see inset of upper panel of Fig. 2a) displays a narrow (pseudo)gap feature of $\sim$50 meV size, in very good agreement with the experimental semiconducting gap of La3Ni2O6 [40]. From the local spectral function, this (pseudo)gap is exclusively linked to Ni-$d_{z^{2}}$, while Ni-$d_{x^{2}-y^{2}}$ is Mott-gapped orbital selectively on a larger scale of $\sim 0.85$ eV. Concerning the local-orbital fillings, the value $n_{x^{2}-y^{2}}=1.10$ marks the expected near-half-filled scenario for Mott-gapped in-plane $d_{x^{2}-y^{2}}$, whereas the higher value $n_{z^{2}}=1.27$ is realized for out-of-plane $d_{z^{2}}$. As shown in the left panel of Fig. 3, the intriguing low-energy spectrum results from a metal-to-insulator transition with lowering temperature. While the Ni-$d_{x^{2}-y^{2}}$ orbital is Mott-gapped already well above 100 K, the Ni-$d_{z^{2}}$ still displays metallic character with a seemingly QP-like peak at $T=150$ K. But in the spectrum at $T=100$ K, the opening of a (pseudo)gap for the latter orbital starts to become visible. Concomitantly, the imaginary part of the Matsubara self-energy $\Sigma(i\omega_{n})$, shown in the right panel of Fig. 3, looks Fermi-liquid like for Ni-$d_{z^{2}}$ over a wide frequency range. Yet it bends downwards for smallest $\omega_{n}$ at $T=50$ K, i.e. signaling the established (pseudo)gap feature. The larger Ni-$d_{x^{2}-y^{2}}$ self-energy also changes qualitative behavior in the (pseudo)-gap regime, pointing towards an intricate coupling scenario between both Ni-$e_{g}$ sectors.
Figures 2b,c display the k-resolved spectral function for La3Ni2O6. The original Ni-$t_{2g}$ DFT flat-band is repelled from the Fermi level by correlations. Instead, the Ni-$d_{z^{2}}$ flat-band is shifted to lowest energy. This correlated Ni-$d_{z^{2}}$ flat band becomes rather incoherent and splits, giving rise to the small (pseudo)gap at low energy. Hence, flat-band physics as observed in La3Ni2O7 is also present in La3Ni2O6, yet with a different twist. The substantial Ni-$e_{g}$ orbital anisotropy in the reduced bilayer renders $d_{x^{2}-y^{2}}$ closer to Mottness, whereas $d_{z^{2}}$ is distant from Mott criticality. But due to its low-energy flat-band electrons scattering with Mott-localized $d_{x^{2}-y^{2}}$ electrons, it still becomes (pseudo)gapped. The question about an intuitive explanation for this Ni-$d_{z^{2}}$ (pseudo)gapping arises. It should be based on the competition between residual (flat-band) hopping and local-moment formation. Because of the already localized Ni-$d_{x^{2}-y^{2}}$ electron with $S=1/2$, the system may gain exchange energy from hybridizing with flat-band Ni-$d_{z^{2}}$ electrons. It has in fact been shown [62], that such a two-band orbital-selective regime is described by a ferromagnetic Kondo-lattice Hamiltonian. Non-Fermi-liquid behavior and pseudogap features are possible solutions of such ferromagnetic/underscreened Kondo-lattice problems [62, 63, 64, 65]. Here, the scattering between localized Ni-$d_{x^{2}-y^{2}}$ and flat-band Ni-$d_{z^{2}}$ leads to a strong (pseudo)gap.
Figure 5: Spin susceptibility and superconducting gap function for the lowest positive Matsubara frequency of the Ni-$d_{z^{2}}$-orbital component in the first Brillouin zone. Note that the even susceptibility in (a) is an order of magnitude smaller than the odd susceptibility in (b). The real part of the intra- and interlayer superconducting gap functions are shown in (c) and (d), respectively. The susceptibility is given in units of states/eV, and the superconducting gaps are in units of meV.
Hole doping.— The behavior with doping now becomes of immediate interest. For instance, hole doping should lower the Ni(3d)(3d) electron count, eventually bringing it in line with the effective $d^{8}$ filling of La3Ni2O7. This could yield a novel correlated-electron setting for unconventional superconductivity. Note that the application of pressure to stoichiometric La3Ni2O6 has already been performed up to 25 GPa, resulting in metallicity but not superconductivity [40]. This is reasonable, since pressure may close the (pseudo)gap because of increasing hybridization. But this alone will likely not be sufficient to enable superconductivity, since additionally also the filling has to be adjusted properly. While in high-pressure La3Ni2O7, the flat band is nearly fully occupied [20], for present stoichiometric La3Ni2O6 the flat-band part seems not too far off half filling. Even without pressure, hole doping alone will abolish the (pseudo)gapped state by effectively raising the kinetic-energy gain above the exchange-energy gain for some doping level.
By employing the virtual-crystal approximation (VCA) for the doped La sites by Sr, we performed DFT+sicDMFT calculations at hole doping levels $x$, and the results are depicted in Fig. 4. From the total spectral function in Fig. 4a, the (quasi-)insulating state is robust up to a significant replacement of about 15% of La by Sr. However, for $x\gtrsim 0.2$ a strong QP resonance appears at the Fermi level. It is of dominant Ni-$d_{z^{2}}$ character as observed from the Ni-$e_{g}$-resolved local spectral functions in Fig. 4c. On the other hand, the Ni-$d_{x^{2}-y^{2}}$ electrons remain essentially localized throughout the studied doping range. While the orbital-selective Mott state is robust with hole doping, the flat band is “released” from its (pseudo)gapped state with increasing $x$. The emerging QP resonance is truly associated with the original flat-band state as displayed in Fig. 4b. There we plot the Bloch-resolved spectral function $A_{\nu}(\omega)$, i.e. the $\bf k$}-integrated interacting spectral function in the Bloch basis $\{\nu\}$, for the most susceptible $\nu=\nu_{\rm f}$. One easily recognizes the large doping-dependent low-energy response. On the other hand, already for interacting states of Bloch label $\nu_{\rm f}\pm 1$, there is only weak low-energy spectral change with doping visible (grey lines in Fig. 4b). This renders obvious that the key correlation effects are local in Bloch space $\{\nu\}$, and with hole doping one unlocks the dominant low-energy flat band dispersion from its (quasi-)insulating nature. This (still weak) Ni-$d_{z^{2}}$ dispersion is finally seen in the k-resolved spectral function of Fig. 4d for $x=0.225$, leading to electron pockets at X and hole pockets at M in the interacting Fermi surface, as shown in Fig. 4e.
Superconducting instability.— It is tempting to consider potential Cooper-pairing due to spin fluctuations in the intriguing Ni-$d_{z^{2}}$ flat-band regime for hole doping $x=0.225$. A phase transition to the superconducting state requires numerical evaluation of the pairing vertex, which consists of the spin and charge susceptibilities and related irreducible vertices. We follow the prescription developed previously for multiorbital systems [66], and applied to La3Ni2O7 [35, 67], where the effective vertices are independent of frequency and momentum. But here, the Green’s functions are taken from the converged DFT+sicDMFT calculation. For further details, we refer to the Supplemental Material.
The spin and charge fluctuations are calculated using the Random Phase Approximation (RPA) for both Ni-$e_{g}$ orbitals. However, the Ni-$d_{x^{2}-y^{2}}$ contributions are calculated to be negligible, leaving the stage for Ni-$d_{z^{2}}$. In bilayer systems with only on-site interactions, complete information requires calculating even and odd spin/charge susceptibilities $\chi^{\rm even/odd}=\chi_{\parallel}\pm\chi_{\perp}$ [68, 34]. The former are shown in Figs. 5a,b, respectively, for an effective interaction corresponding to a Stoner enhancement factor of $\alpha_{\rm sp}=0.9$. A magnetic instability is indicated by $\alpha_{\rm sp}=1$. Note that the spin fluctuations in the odd channel are an order of magnitude larger than in the even channel. The corresponding magnetic instability would thus have antiferromagnetic interlayer intrabilayer order.
The leading superconducting solution is a $s_{\pm}$-wave with large interlayer component, similar to what is frequently discussed for La3Ni2O7. The real part of the gap structure of the leading instability is visualized in Figs. 5c,d, focusing on the lowest Matsubara frequency Ni-$d_{z^{2}}$ components. The imaginary part has a similar shape and is smaller in magnitude. The shown $s_{\pm}$ solution is strongly dominant for all reasonable effective interactions and already about unity for $\alpha_{\rm sp}=0.9$, indicating a transition to the superconducting state. Note that the interlayer component is not affected by on-site Coulomb repulsion, such that no sign changes are required to avoid the repulsion. Furthermore, the combinations $\Delta_{\parallel}\pm\Delta_{\perp}$ have opposite signs, which is also expected for the $s_{\pm}$ scenario in La3Ni2O7 [34]. For the intralayer part, it is important to mention that there is a non-trivial frequency dependence, with the higher-frequency components being dominated by the opposite sign likely protecting the solution against Coulomb repulsion (see Fig. S1 in the Supplemental Material).
For comparison, we also solved the frequency-dependent gap equation for pressurized La3Ni2O7 using the DFT+sicDMFT Green’s functions from Ref. [20] (see Supplemental Material for details). There, the magnetic instability for the correlated state is dominated by Ni-$d_{x^{2}-y^{2}}$. Different competing gap structures of multiorbital and multilayer type, including not only a $s_{\pm}$-wave but also $d_{x^{2}-y^{2}}$ gap structure, have comparable eigenvalues. The eigenvalues approach unity only in close proximity to the magnetic instability. Therefore, from the correlated DFT+sicDMFT picture, hole-doped La3Ni2O6 may even offer a better platform to realize strong-coupling unconventional superconductivity of $s_{\pm}$ type with a dominant interlayer Ni-$d_{z^{2}}$ component than La3Ni2O7. At the same time, one has to bear in mind that other density-wave like instabilities, not included in the present consideration, might affect this conclusion.
Summary.— Using advanced first-principles many-body theory we identified the La3Ni2O6 compound as a novel striking correlated material in the family of transition-metal oxides. The reduced RP bilayer compound serves as a unique example where orbital-selective Mott and flat-band physics join in to form a correlated (quasi-)insulator. It hosts orbital-selective Mott-localized Ni-$d_{x^{2}-y^{2}}$ electrons and a flat-band electron state of Ni-$d_{z^{2}}$ character which is gapped by $\sim$50 meV. This gapping originates from scattering between both orbital sectors within an anomalous Kondo-insulator scenario. The reduced bilayer system thus serves as the long-sought explicit materials connection between heavy-fermion and correlated-oxide physics. Doping the flat-band state offers an interesting route for strong-coupling unconventional superconductivity with a dominant $s_{\pm}$-wave symmetry solution residing on Ni-$d_{z^{2}}$-bilayer splitted orbitals. This finding links the already established superconducting regimes of RP- and infinite-layer-based nickelates. Last but not least, after the recent discovery of superconducting La3Ni2O7 under compressive strain [69, 70], the present work paves the way for ambient-pressure high-$T_{\rm c}$ superconductivity in bulk nickelates.
I Acknowledgements
The work is supported by the German Research Foundation within the bilateral NSFC-DFG Project ER 463/14-1. Computations were performed at the Ruhr-University Bochum and the JUWELS Cluster of the Jülich Supercomputing Centre (JSC) under project miqs.